OpenStudy (anonymous):

If an airplane is flying directly north at 300.0 km/h, and a crosswind is hitting the airplane at 50.0 km/h from the east, what is the airplane's resultant velocity?

OpenStudy (anonymous):

\[300^2+50^2=v^2\]

OpenStudy (ash2326):

@joe4321 See the figure the wind will change the orientation of the plane and hence it's velocity

OpenStudy (ash2326):

 Now the plane will move in this direction its velocity V will be given as \[V^2=300^2+50^2\] \[V=\sqrt{300^2+50^2} km/h\] do you understand this?